OpenStudy (anonymous):

A horizontal force of 160 N is exerted on a 50.0 kg bed and the bed is moved 5.0 m horizontally in 20 seconds. What is the work done by the force? What is the power? Show formula and work for both calculations.

OpenStudy (ash2326):

F=160 N M=50 kg Work done = Force*Displacement Wok Done=160*5=800 Joules Power=Work/ Time Power=800/20 =40 watts